Can you please help me solve using the substitution method?
x + y = 1
x + y = 2

Answer by checkley71(8403)

X+Y=1
X=1-Y
(1-Y)+Y=2
1-Y+Y=2
0Y=1 NO SOLUTION
------------------------------------------
X+Y=1 & X+Y=2 CANNOT BOTH BE VALID BECAUSE 1 DOES NOT EQUAL 2
PROOF
SET (X+Y)=Z
NOW SOLVE
Z=1 & Z=2
Z CANNOT BE 1 & 2 IN THE SAME SET OF LINEAR EQUATIONS.